stock has 8% annual dividend and $100 par value and was sold a t$99.50 per share. Flotation cost of $1.50 per share.

A) Calculate cost of preferred stock

B) If the firm sells the preferred stock with a 10% annual dividend and nets $90.00 after flotation cost, what is its cost?
Answer #1

1.

Annual dividend=100*8%=8

Hence cost of preferred stock=Annual dividend/(Current price-Flotation cost)

=8/(99.5-1.5)

=8/98

which is equal to

=8.16%(Approx).

2.Annual dividend=100*10%=10

Hence cost of preferred stock=Annual dividend/Net price

=10/90

=11.11%(Approx).
Answer #2

A) To calculate the cost of preferred stock, we can use the formula:

Cost of preferred stock = (Dividend per share / Net issuance price) + Flotation cost per share

Given that the annual dividend is 8% of the $100 par value and the stock is sold at $99.50 per share with a flotation cost of $1.50 per share, we can calculate the cost of preferred stock:

Dividend per share = 8% of $100 par value = 0.08 * $100 = $8.00 Net issuance price = Selling price - Flotation cost = $99.50 - $1.50 = $98.00

Cost of preferred stock = ($8.00 / $98.00) + $1.50 Cost of preferred stock ≈ 0.0816 + $1.50 ≈ $1.5816

B) If the firm sells the preferred stock with a 10% annual dividend and nets $90.00 after flotation cost, we can calculate the cost of the preferred stock using the same formula:

Dividend per share = 10% of $100 par value = 0.10 * $100 = $10.00 Net issuance price = $90.00

Cost of preferred stock = ($10.00 / $90.00) + $1.50 Cost of preferred stock ≈ 0.1111 + $1.50 ≈ $1.6111

So, the cost of preferred stock for the firm, in this case, is approximately $1.6111 per share.


answered by: mervetokaz